i need to know how to do somethings such as i have an 85% alcohol solution and a 50%alcohol solution and i need a 65% solution. and i need 55 liters all together. how many liters of each do i need? i have the eqautions
x=how many liters of 85%
y=how many liters of 50%
i have x+y=55 (liters)
and i have .85x+.50y= .65
please reply as soon as possible, thank you

Answer by stanbon(75887)

x=how many liters of 85%
y=how many liters of 50%
--------------------------------
i have x+y=55 (liters)
and i have .85x+.50y= .65*55
------------------------------
Multiply thru the 1st by 85 and the 2nd by 100 to get:
85x + 85y = 85*55
85x + 50y = 65*55
-----------------------
Subtract the 2nd from the 1st and solve for "y":
35y = 20*55
y = 31.43 liters (amt. of 50% solution needed in the mixture)
Since x + y = 55, x = 55-31.43 = 23.57 liters (amt of 85% needed in the mix)
===========================================================
Cheers,
Stan H.